OpenStudy (anonymous):

Two rainstorms occurred in one week in a certain area. In the first storm, 30ml of rain fell per hour, and in the second storm, 25ml of rain fell per hour. Rain lasted for a total of 35 hours, resulting in a total rainfall of 975ml. How long was each storm?

OpenStudy (anonymous):

messed that up didn't i put x = hours at 30, 35-x = hours at 25 and solve \[30x+25(35-x)=975\]

OpenStudy (anonymous):

\[35x+875-25x=975\] \[5x+875=975\] \[5x=100\] \[x=20\]

OpenStudy (anonymous):

answer is 20 at 30 ml per hour and therefore 15 at 25ml per hour